# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ Appendix B Proofs of Theorems, Corollaries and Propositions
$\begin{split}&\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,n/(2K)}(\mathfrak{F})+\delta\right]\\ &\qquad\qquad\qquad\qquad\qquad\geq 1-2\exp(-n\delta^{2}/Kb^{2})-2(n/2K-1)\beta(K)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad>1-2\exp[-n\delta^{2}/(4Kb^{2})]-2(n/2K)\beta(K).\end{split}$ (B.8)
Notice that, in order to ensure that both $2\exp(-n\delta^{2}/4Kb^{2})$ and $2(n/2K)\beta(K)$ vanish to zero (under Assumption (VII) for $\beta(K)$), it is tempting to apply (B.8) with $K=n^{\alpha}$ for some $\alpha<1$. Unfortunately, there is no reason for such a $K$ to be an integer. A solution, would be to let $K=\lfloor n^{\alpha}\rfloor$, but in this case $n/(2K)$ might not be an integer. To address such issues, it is necessary to consider a careful modification of (B.8). To this end write the Euclidean division $n=n^{\prime}+r$ where $n^{\prime}=2Kh\leq n$, $h=\lfloor n/(2K)\rfloor$ and $0\leq r<2K$. Then, under the common marginal assumption and recalling also Proposition 1 in Mohri & Rostamizadeh (2008) together with the triangle inequality and the fact that the functions $f$ within $\mathfrak{F}$ are $b$–uniformly bounded, we have that
@@
where the last inequality follows directly from the definition of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)$ together with the fact that $0\leq r<2K$. Applying now (B.8) to $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)+4bK/n$ yields
@@
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\delta\right]\geq 1-2\exp(-s_{n}\delta^{2}/2b^{2})-2s_{n}\beta(\lfloor\sqrt{n}\rfloor)$
@@
Proof of Lemma 6.2.
The proof follows the arguments used in Chapter 2 of Doukhan (1994) to study general Markov chains. In particular, when $(x_{t})_{t\in\mathbb{Z}}$ is a stationary Markov chain with invariant distribution $\pi(\cdot)$ and transition kernel $P(\cdot,\cdot)$, a result proven in Davydov (1974) and recalled in page 87–88 of Doukhan (1994) gives
$\beta(k)=\mathbb{E}_{x\sim\pi}\|P^{k}(x,\cdot)-\pi(\cdot)\|_{\textsc{TV}}.$
When $-1<\theta<1$, standard results for the ar(1) model in Lemma 6.2 lead to the invariant distribution $\pi=\mbox{N}(\psi/(1-\theta),\sigma^{2}/(1-\theta^{2}))$.
As for $P^{k}(x,\cdot)$, notice that, under such an ar(1) model, we can write
@@
Therefore, $P^{k}(x,\cdot)=\mbox{N}(\theta^{k}x+\psi\sum_{l=0}^{k-1}\theta^{l},\sigma^{2}\sum_{l=0}^{k-1}\theta^{2l})$. Moreover, notice that, by direct application of standard properties of finite power series, we have $\sum_{l=0}^{k-1}\theta^{l}=(1-\theta^{k})/(1-\theta)$ and $\sum_{l=0}^{k-1}\theta^{2l}=\smash{(1-\theta^{2k})/(1-\theta^{2})}$. Since our goal is to derive an upper bound for $\beta(k)$ and provided that the kl divergence among Gaussian densities is available in closed form, let us first consider the Pinsker’s inequality
@@
$D_{\textrm{KL}}(P^{k}(x,\cdot),\pi(\cdot))$
@@
Proof of Proposition 6.1.
Under Assumption (VII), for any fixed $\delta>0$, we have that $\sum_{n>0}[2\exp(-s_{n}\delta^{2}/2b^{2})+2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})]<\infty$. Therefore, combining Lemma 6.1 with Assumption (IV), both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ converge to 0 almost surely as $n\to\infty$, by the Borel–Cantelli Lemma. As a result, since $-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})-\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*}),$ it holds that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$. To conclude it suffices to apply again the proof of Theorem 1 in Jiang et al. (2018); see also proof of Corollary 3.1. ∎
Proof of Theorem 6.1.
To prove Theorem 6.1 we will follow the same line of reasoning as in the proof of Theorem 3.2. However, in this case we leverage Lemma 6.1 instead of Lemma 2.1. To this end, letting $\delta=c_{1}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n}$ with $c_{1}\geq 2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$ and $\mathfrak{R}_{s_{n}}=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,s_{n}}$, we obtain, under Assumption (VII), Equation (B.9) below, instead of (B.3).
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq c_{1}\right]\\ \geq 1-2\exp[-s_{n}(c_{1}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi}).$ (B.9)
As in Theorem 3.2, let $c_{1}=\bar{\varepsilon}_{n}/3$ and notice that, by the settings of Theorem 6.1, for $n$ large enough $\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$. Therefore, applying (B.9) to $y_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3$, leads to the following upper bound
@@
Recall that, under the settings of Theorem 6.1, we have that $\sqrt{n}\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{s_{n}}(\mathfrak{F})\to\infty$ and, therefore, $s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}\sim s_{n}\bar{\varepsilon}^{2}_{n}/9\sim\sqrt{n}\bar{\varepsilon}^{2}_{n}\to\infty$. Combining this result with the fact that $2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})\to 0$ as $n\to\infty$ (recall (VII)), it follows that the right end side of the above equation goes to $1$ as $n\to\infty$. Hence, in the remaining part of the proof, we will restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$.
@@
$\begin{split}&\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\left(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\}\right)\\ &\qquad=\frac{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}.\end{split}$ (B.10)
To upper bound the ratio in (B.10), let us first derive an upper bound for the numerator. To this end, consider the triangle inequality $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ (recall that $\mathcal{D}_{\mathfrak{F}}$ is a semimetric), along with the previously–proved result that the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$ has $\mathbb{P}_{y_{1:n}}$–probability going to $1$. Therefore, for large $n$, we have
@@
$\qquad\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta}){+}\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}/3{+}c_{2}\right]\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right].$
Now, note that $\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right]=\int_{\theta\in\Theta}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta\right]\pi({\rm d}\theta)$. Therefore, by applying (B.9) to $z_{1:n}$ yields,
@@
This controls the numerator in (B.10). As for the denominator, defining the event $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$ and applying again the triangle inequality, we have that
@@
The last inequality follows from that fact that we can restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$, and that we are integrating over $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$. Let us now apply again equation (B.9) to $z_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3$, to further lower bound the last term of the above inequality by
@@
Note that, by Assumption (II), $\pi(\mbox{E}_{n})\geq c_{\pi}(\bar{\varepsilon}_{n}/3)^{L}$. Moreover, as shown before, the quantity $1-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})$ goes to $1$ when $n\to\infty$, which also implies that, for a large enough $n$, $21-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})>1/$ 2. Therefore, leveraging both results, the denominator in (B.10) can be lower bounded by $(c_{\pi}/2)\left(\bar{\varepsilon}_{n}/{3}\right)^{L}$. To proceed with the proof, let us combine the upper and lower bounds derived, respectively, for the numerator and the denominator of the ratio in (B.10). This yields, for any integer $K$,
[…]
@@
Replacing $c_{2}$ in (B.11) with $2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\sqrt{(2b^{2}/s_{n})\log(n/\bar{\varepsilon}_{n}^{L})}+4b/\sqrt{n}$ , gives
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\left[\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\right]^{1/2}}\Bigr{\}}\Bigr{)}\\ \leq\frac{4\cdot 3^{L}}{nc_{\pi}}\left[1+\frac{ns_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})}{\bar{\varepsilon}_{n}^{L}}\right].$
To conclude, note that $ns_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})/\bar{\varepsilon}_{n}^{L}=n^{1+L/2}s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})/((n\bar{\varepsilon}_{n}^{2})^{L/2})$ where the numerator goes to 00 and the denominator goes to $\infty$ when $n\to\infty$, under the setting of Theorem 6.1. Therefore, for $n$ large enough,
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\left[\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\right]^{1/2}}\Bigr{\}}\Bigr{)}\\ \leq\frac{4\cdot 3^{L}}{nc_{\pi}},$

### v5

@@ Appendix D Proofs of Theorems, Corollaries and Propositions
$\begin{split}&\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,n/(2K)}(\mathfrak{F})+\delta\right]\\ &\qquad\qquad\qquad\qquad\qquad\geq 1-2\exp(-n\delta^{2}/Kb^{2})-2(n/2K-1)\beta(K)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\geq 1-2\exp[-n\delta^{2}/(4Kb^{2})]-2(n/2K)\beta(K).\end{split}$ (D.8)
Notice that, in order to ensure that both $2\exp(-n\delta^{2}/4Kb^{2})$ and $2(n/2K)\beta(K)$ vanish to zero (under Assumption (VII) for $\beta(K)$), it is tempting to apply (D.8) with $K=n^{\alpha}$ for some $10<\alpha<$ 1. Unfortunately, there is no reason for such a $K$ to be an integer. A solution, would be to let $K=\lfloor n^{\alpha}\rfloor$, but in this case $n/(2K)$ might not be an integer. To address such issues, it is necessary to consider a careful modification of (D.8). To this end write the Euclidean division $n=n^{\prime}+r$ where $n^{\prime}=2Kh\leq n$, $h=\lfloor n/(2K)\rfloor$ and $0\leq r<2K$. Then, under the common marginal assumption and recalling also Proposition 1 in Mohri and Rostamizadeh (2008) together with the triangle inequality and the fact that the functions $f$ within $\mathfrak{F}$ are $b$-uniformly bounded, we have that
@@
where the last inequality follows directly from the definition of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)$ together with the fact that $0\leq r<2K$. Applying now (D.8) to $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)$ yields
@@
$\mathbb{P}_{x_{1:n}}\Bigl{[}\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\delta\Bigr{]}\geq 1-2\exp(-s_{n}\delta^{2}/2b^{2})-2s_{n}\beta(\lfloor\sqrt{n}\rfloor)$
@@
Proof of Lemma C.5.
The proof follows the arguments used in Chapter 2 of Doukhan (1994) to study general Markov chains. In particular, when $(x_{t})_{t\in\mathbb{Z}}$ is a stationary Markov chain with invariant distribution $\pi(\cdot)$ and transition kernel $P(\cdot,\cdot)$, a result proven in Davydov (1974) and recalled in page 87–88 of Doukhan (1994) gives $\beta(k)=\mathbb{E}_{x\sim\pi}\|P^{k}(x,\cdot)-\pi(\cdot)\|_{\textsc{TV}}.$ When $-1<\theta<1$, standard results for the ar(1) model in Lemma C.5 lead to the invariant distribution $\pi=\mbox{N}(\psi/(1-\theta),\sigma^{2}/(1-\theta^{2}))$.
As for $P^{k}(x,\cdot)$, notice that, under such an ar(1) model with starting point $x$, we can write
@@
Therefore, $P^{k}(x,\cdot)=\mbox{N}(\theta^{k}x+\psi\sum_{l=0}^{k-1}\theta^{l},\sigma^{2}\sum_{l=0}^{k-1}\theta^{2l})$.
Moreover, notice that, by direct application of standard properties of finite power series, we have $\sum_{l=0}^{k-1}\theta^{l}=(1-\theta^{k})/(1-\theta)$ and $\sum_{l=0}^{k-1}\theta^{2l}=\smash{(1-\theta^{2k})/(1-\theta^{2})}$. Since our goal is to derive an upper bound for $\beta(k)$ and provided that the kl divergence among Gaussian densities is available in closed form, let us first consider the Pinsker’s inequality
@@
$\mathcal{D}_{\textrm{KL}}(P^{k}(x,\cdot),\pi(\cdot))$
@@
Proof of Proposition C.6.
Under Assumption (VII), for any fixed $\delta>0$, we have that $\sum\nolimits_{n>0}[2\exp(-s_{n}\delta^{2}/2b^{2})+2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})]<\infty.$ Therefore, combining Lemma C.2 with Assumption (IV), both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ converge to 0 almost surely as $n\to\infty$, by the Borel–Cantelli Lemma. As a result, since
[…]
it holds that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$. To conclude it suffices to apply again the proof of Theorem 1 in Jiang, Wu and Wong (2018); see also the proof of Corollary 3.2. ∎
Proof of Theorem C.7.
To prove Theorem C.7 we will follow the same line of reasoning as in the proof of Theorem 3.3. However, in this case we leverage Lemma C.2 instead of Lemma 2.6. To this end, letting $\delta=c_{1}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n}$ with $c_{1}\geq 2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$ and $\mathfrak{R}_{s_{n}}=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,s_{n}}$, we obtain, under Assumption (VII), Equation (D.9) below, instead of (D.3).
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq c_{1}\right]\\ \geq 1-2\exp[-s_{n}(c_{1}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi}).$ (D.9)
As in Theorem 3.3, let $c_{1}=\bar{\varepsilon}_{n}/3$ and notice that, by the settings of Theorem C.7, for $n$ large enough $\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$. Therefore, applying (D.9) to $y_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3$, leads to the following upper bound
@@
Recall that, under the settings of Theorem C.7, we have that $\sqrt{n}\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{s_{n}}(\mathfrak{F})\to\infty$ and, therefore,
$s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}\sim s_{n}\bar{\varepsilon}^{2}_{n}/9\sim\sqrt{n}\bar{\varepsilon}^{2}_{n}\to\infty.$
Combining this result with the fact that $2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})\to 0$ as $n\to\infty$, it follows that the lower bound for $\mathbb{P}_{y_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\right]$ goes to $1$ as $n\to\infty$. Hence, in the remaining part of the proof, we will restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$.
@@
$\begin{split}&\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\left(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\}\right)\\ &\qquad=\frac{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}.\end{split}$ (D.10)
To upper bound the ratio in (D.10), let us first derive an upper bound for the numerator. To this end, consider the triangle inequality $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ (recall that $\mathcal{D}_{\mathfrak{F}}$ is a semimetric), along with the previously-proved result that the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$ has $\mathbb{P}_{y_{1:n}}$–probability going to $1$. Hence, for $n$ large enough we have
@@
$\qquad\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}/3+c_{2}\right]\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right],$
where $\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right]=\int_{\theta\in\Theta}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta\right]\pi({\rm d}\theta).$
Therefore, leveraging the above result and applying (D.9) to $z_{1:n}$ yields,
@@
This controls the numerator in (D.10). As for the denominator, defining the event $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$ and applying again the triangle inequality, we have that
@@
The last inequality follows from that fact that we can restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$, and that we are integrating over $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$. Let us now apply again (D.9) to $z_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3$, to further lower bound the last term of the above inequality by
@@
Note that, by Assumption (II), $\pi(\mbox{E}_{n})\geq c_{\pi}(\bar{\varepsilon}_{n}/3)^{L}$. Moreover, as shown before, the quantity $1-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})$ goes to $1$ when $n\to\infty$, which also implies that, for a large enough $n$,
$21-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})>1/$ 2 .
Therefore, leveraging both results, the denominator in (D.10) can be lower bounded by the term $(c_{\pi}/2)\left(\bar{\varepsilon}_{n}/{3}\right)^{L}$.
To proceed with the proof, let us combine the upper and lower bounds derived, respectively, for the numerator and the denominator of the ratio in (D.10). This yields, for any integer $K$,
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\})\\ \leq\frac{2\exp[-s_{n}(c_{2}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]+2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})}{(c_{\pi}/2)(\bar{\varepsilon}_{n}/3)^{L}},$ (D.11)
@@
Replacing $c_{2}$ in (D.11) with $2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\sqrt{(2b^{2}/s_{n})\log(n/\bar{\varepsilon}_{n}^{L})}+4b/\sqrt{n}$ , gives
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\Bigl{(}\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}}\Bigr{\}}\Bigr{)}\\ \leq\frac{4\cdot 3^{L}}{nc_{\pi}}\Bigl{(}1+\frac{ns_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}.$
To conclude, note that
$ns_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})/\bar{\varepsilon}_{n}^{L}=n^{1+L/2}s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})/((n\bar{\varepsilon}_{n}^{2})^{L/2}),$
where the numerator goes to 00 and the denominator goes to $\infty$ when $n\to\infty$, under the setting of Theorem C.7. Therefore, with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\to\infty$, we have
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\Bigl{(}\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}}\Bigr{\}}\Bigr{)}\\ \leq\frac{4\cdot 3^{L}}{nc_{\pi}},$